A 2 tone car is towed by a 2.5 tone truck which moves from rest and reaches a speed o 2m/s before tightens on the stationary car. The rope then stretches taking up the strain fro 0.15 seconds, find the common speed when the rope ceases to stretch


1
Expert's answer
2020-02-12T10:41:07-0500
"Mu=(m+M)v"

"(2.5)(2)=(2+2.5)v"

"v=1.1\frac{m}{s}"